# CP violation in two-meson Tau decays induced by heavy new physics

### v1

@@ Abstract
We apply the effective field theory formalism that was used to study CP violation induced by heavy new physics in the $\tau\to K_{S}\pi\nu_{\tau}$ decays to the other two-meson tau decay channels. We focus on the rate and the forward-backward asymmetries, that are predicted using current bounds on the complex Wilson coefficients of the effective Lagrangian. We discuss our outcomes for the modes with $(K/\pi)\pi^{0}$ and $KK_{S}$, that can be studied at Belle-II and a super-tau-charm facility. Our main finding is that current and forthcoming experiments would be sensitive to the maximum allowed CP rate asymmetry in the $K^{\pm}K_{S}$ modes if a precision of $5\%$ is reached on this observable. For the $\pi^{\pm}\pi^{0}$ channels, new physics would be best probed through the forward-backward asymmetry. Disentangling new sources of CP violation would be most challenging in $K^{\pm}\pi^{0}$ and the other modes.
@@ 1 Introduction
Violation of the invariance under the combined symmetry operations of charge conjugation (C) and parity (P), CP, is a required ingredient -within any Lorentz invariant local quantum field theory with a Hermitian Hamiltonian- to understand the baryon asymmetry of the Universe, as pointed out by Zakharov in his seminal paper [1]. It also needs baryon number violation (which could be triggered violating initially the lepton number [2]), non-conservation of C, and interactions out of thermal equilibrium. In principle all these conditions can be met within the Standard Model (SM) [3, 4, 5], though it fails to explain the current matter-antimatter asymmetry [6]. This conundrum enhances the case for searches of additional sources of CP violation, that can eventually solve this puzzle.
@@
In the quark sector, CP violation in the SM arises only from an irreducible phase in the Cabibbo-Kobayashi-Maskawa (CKM) mixing matrix [7, 8], and has been measured in $K$, $B_{(s)}$ and $D$ mesons [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. To date, no data from hadron decays or meson mixing seems to require CP violation beyond the SM [6]. Conversely, CP violation in the lepton sector has not been discovered yet, although there are hints for it both in neutrino oscillation measurements by Nova [20] and T2K [21], and in global fits [6]. If confirmed, it would remain to be seen if it can be explained by the single CP violating phase that appears in the lepton analogue of the CKM matrix, the Pontecorvo-Maki-Nakagawa-Sakata matrix [22, 23, 24], or new physics is called for.
@@
Hadron tau decays are not only a privileged arena for learning about non-perturbative QCD in a clean environment, but also a very useful tool in the quest for new physics [25, 26]. Regarding CP violation, this was emphasized for the first generation B-factories, BaBar and Belle, and forthcoming facilities [27, 28, 29, 30, 31]. CP non-conservation in semileptonic tau decays was boosted by the anomaly reported by BaBar [32] on the difference between the decay rates of a $\tau^{+}$ and a $\tau^{-}$ into final states with a correspondingly charged pion and a $K_{S}$ (inclusive in possible neutral pions), normalized by their sum. While this CP violating observable should be $(0.36\pm 0.01)\%$ in the Standard Model –due to neutral kaon mixing and accounting for the precise BaBar experimental setting– [33, 34, 35], the measurement came out with opposite sign, $-(0.36\pm 0.23\pm 0.11)\%$ (with statistical and systematic errors in turn), being $2.8$ standard deviations away from the theory prediction. Although there were claims that this effect could be due to heavy new physics (particularly of tensor type [36, 37, 38]), the key equality of the tensor and vector form factors phases in the elastic region allowed to disfavor this explanation unless an unnatural fine-tuning was involved [39] (see also refs. [40, 41, 42, 43, 44] [^1] No CP violation was found in the binned analysis of these decays by Belle [45]. However, its precision was not enough to clarify the BaBar anomaly (neither in the pioneering CLEO studies [46, 47])..
@@
The aim of this paper is to extend the work done in the previous references to the other two-meson tau decays of interest ($\pi\pi$, $K\bar{K}$ and $K\pi^{0}$ channels  [^2] The $\pi\eta^{(\prime)}$ channels have not been measured yet, and the hadronic uncertainty in them is big [48, 49], so any current prediction would be subject to large uncertainties. CP conserving new physics imprints in these decays were the subject of ref. [50]. We will comment briefly on the $K(\eta/\eta^{\prime})$ channels, also subject to large uncertainties, at the end of section 5.  [^3] Model-dependent studies have been carried out in e. g. refs. [51, 52, 53, 54, 55, 56].  [^4] We will not address in this work the case of polarized taus, which would allow further tests of $CP$ and time-reversal symmetry violations, along the lines of ref. [57].). In the same way that heavy new physics contributions to the CP violating $\tau\to K_{S}\pi\nu_{\tau}$ observables is very much restricted, model-independently, upon translating the constraints on the SM effective-field theory (EFT) [58, 59] operators to the low-energy EFT (LEFT) ones, it will happen similarly in the other two-meson tau decay modes. Our purpose is to compute the maximum amount of CP violation that can be due to heavy new physics in these processes. In this way, any observation beyond our predicted upper limits  [^5] Within the Standard Model CP violation is negligibly small in all tau decays without a neutral Kaon, see for instance ref. [60]. will either be due to light new physics or signal underestimated uncertainties, which would deserve a close scrutiny. The discussed physics shall be of interest for Belle-II [61] and future super-tau-charm factory [62, 63, 64] searches.
@@ 2 Low-energy Effective Field Theory Computation
The most general Fermi-like effective Lagrangian for the $\tau^{-}\to\bar{u}D\nu_{\tau}\,(D=d,s)$ decays is given by   [^6] In our conventions, $\sigma^{\mu\nu}=\frac{i}{2}[\gamma^{\mu},\gamma^{\nu}]$ and $\epsilon^{0123}=+1$.[65, 50, 39]
@@
The other decay channels of interest are trivially obtained from the $K^{-}\pi^{0}$ one, replacing $p_{K}$ ($p_{\pi}$) by the charged (neutral) meson momentum and substituting the $1/2$ factor in eq. (2) by $1/\sqrt{2}$ for the $K^{-}K^{0}$ and $\bar{K}^{0}\pi^{-}$ modes and by $1$ in the $\pi^{-}\pi^{0}$ channel. The $K^{-}\eta^{(\prime)}$ channels are a bit more complicated. Relative to the $K^{-}\pi^{0}$ modes, one needs to substitute $1\to c^{V}_{K^{-}\eta^{(\prime)}}$ in the $F_{+}(s)$ term and $1\to-c^{S}_{K^{-}\eta^{(\prime)}}$ in the $F_{0}(s)$ term, where $c^{V}_{K^{-}\eta^{(\prime)}}=-\sqrt{\frac{3}{2}}$, $c^{S}_{K^{-}\eta}=-\frac{1}{\sqrt{6}}$ and $c^{S}_{K^{-}\eta^{\prime}}=\frac{2}{\sqrt{3}}$ [66]. The scalar matrix element, $H$, demands to modify $1\to c^{S}_{K^{-}\eta^{(\prime)}}$, in order to obtain the $K^{-}\eta^{(\prime)}$ result from the $K^{-}\pi^{0}$ expression. For the tensor matrix element $H^{\mu\nu}$, the $K^{-}\eta^{(\prime)}$ result is found from the $K^{-}\pi^{0}$ one by changing $1\to c^{V}_{K^{-}\eta^{(\prime)}}$.
The observables of interest are quoted in the following. All of them check the results in ref. [44] accounting for the relative factor of $1/2$ entering our results for the (differential) decay width:
@@
where $\lambda(a,b,c)=a^{2}+b^{2}+c^{2}-2(ab+ac+bc)$ and $S_{EW}\sim 1.02$ [70] encodes the universal short-distance electroweak radiative corrections (which only affect the SM contribution, but are written here as an overall factor for simplicity, since the induced error of this approximation is subleading  [^8] We are not considering observables inclusive in photon emission, so that the process-dependent long-distance electromagnetic correction [67, 68, 69] is not applied.). Integrating the angular dependence  [^9] We recall that [71] $\cos\alpha=\cos\beta\cos\Psi+\sin\beta\sin\Psi\cos\phi$, where the $\alpha,\beta,\phi,\Psi$ angles are defined in this reference., we get
@@
where $\delta_{+,T}(s)$ denote the phases of the vector and tensor form factors, in turn. The expression in the second and third lines of eq. (11) corresponds to the new physics (NP) contribution to $A_{CP}^{rate}$, which coincides with $A_{CP}^{rate}$ for those modes without neutral Kaons. In the $\tau^{\pm}\to K^{\pm}K_{S}\nu_{\tau}$ case, this observable is dominated by the SM contribution, according to [36]
@@
For completeness, we note that these observables can be made bin-dependent straightforwardly, following ref. [44], namely
@@ 3 Form-factors input
We will use the dispersive form factors that were employed in e. g. ref. [72], which bound the modulus of the Wilson coefficients $\hat{\epsilon}_{S,T}^{D}$ using exclusive semileptonic tau decays into one and two mesons. We will illustrate this hadron input here, taking as an example the simplest di-pion $F_{+}$ case, [73, 74, 75]. The other dispersive form factors in this channel and those for the remaining decay modes were worked out in e. g. refs. [48, 76, 77, 78, 79, 80, 81, 82, 83, 40, 66, 84, 85] and agree well with data, given the present uncertainties. We will briefly comment on the sources of theory uncertainty for $F_{+}^{\pi\pi}$, which is the best known, below.
@@
where $\alpha_{1,2}$ are subtraction constants (the other one is fixed by the conservation of the vector current, so that $F_{+}(0)=1$) giving the slope and curvature of the threshold expansion of $F_{+}(s)$, where Chiral Perturbation Theory provides useful constraints [86, 87, 88, 89]. In practice, since the phaseshift $\delta_{+}(s)$ is not known up to infinity, several approaches are used to account for the associated uncertainty, as detailed in the quoted refs. In the resonance region we obtain the seed for $\tan\delta_{+}(s)=\frac{\Im mF_{+}(s)}{\Re eF_{+}(s)}$ from Resonance Chiral Theory ($R\chi T$[90, 91], with parameters encoding the resonance properties (their pole positions and interferences among them) that are fitted to data. Our form factors comply with the leading asymptotic behaviour predicted by perturbative QCD [92].
@@
In the remainder of this section, we summarize the main sources of theory uncertainty, analyzed minutely in ref. [75] for $F_{+}^{\pi\pi}(s)$. In the elastic zone (which, neglecting the phase-space suppressed four-pion contributions extends up to the di-Kaon channel, with invariant mass around a GeV), the form factor phase is known with extreme precision, thanks to the Roy-Steiner equation analyses of pion scattering data by the Madrid [93] and Bern [94] groups (and to the Watson theorem [95] that ensures the equality of the scattering and form factor phases in this region). Ref. [75], working within the excellent approximation of isospin symmetry, uses these results in the elastic domain, including the possible small variation of the precise energy at which the phase reaches $\pi/2$, signaling the $\rho(770)$ resonance position. Another source of uncertainty corresponds to the matching point between the elastic phase and the one obtained using as seed the $R\chi T$ expression, where values in the interval $[0.9,1]$ GeV are employed. For the $\rho^{\prime(\prime)}$ resonances widths, both the simplest case with only the di-pion channel and also that adding the di-Kaon and $\omega$-$\pi$ cuts were considered. As commented before, there is a lack of experimental information on $\delta_{+}^{\pi\pi}(s)$ above a given energy, and we only know that it must approach $\pi$ asymptotically [96]. Several methods quantified this uncertainty in ref. [75]: cutting the upper limit ($s_{cut}$) of the dispersive integral in $4$, or $9$ GeV2, or keeping it up to $\infty$; modifying the phase for $s\sim s_{cut}$ to avoid spurious singularities; including additional inelastic contributions via a conformal polynomial, etc. It was also verified that the number of subtraction constants was optimal using the reduced $\chi^{2}$ as a criterion. The impact of the $\tau^{-}\to K^{-}K_{S}\nu_{\tau}$ data was also explored, although it is not yet precise enough to improve our knowledge of the $\rho^{\prime}-\rho^{\prime\prime}$ interference.
@@
In this way, ref. [75] quantified the uncertainty of $F_{+}^{\pi\pi}(s)$, whose modulus and phase are represented in figs. 1 and 2. Following ref. [39], we estimated the difference between the vector and tensor form factors in the inelastic region  [^10] They coincide in the elastic one, as elegantly explained in ref. [39]. ($inel$) using $|\delta_{+}(s)-\delta_{T}(s)|=2\delta_{+}^{inel}(s)$. The tensor form factor phase, $\delta_{T}(s)$, thus obtained is compared in fig. 3 to $\delta_{+}(s)$. Fig. 4 plots $|F_{T}(s)|/F_{T}(0)$. In this case, we lack information on the slope parameters, so we use a once-subtracted dispersion relation, which increases the sensitivity to $s_{cut}$. As a result, the uncertainty on $|F_{T}(s)|$ is larger than for $|F_{+}(s)|$, despite being their phases identical in the elastic regime.
@@
Figure 1: Pion vector form factor modulus as a function of the $\pi\pi$ invariant mass. The solid curve represents the central value, and the dashed lines cover the one sigma uncertainties, according to ref. [75].
@@
Figure 2: Pion vector form factor phase as a function of the $\pi\pi$ invariant mass. The solid curve represents the central value, and the dashed lines cover the one sigma uncertainties, according to ref. [75].
@@
Figure 3: Pion tensor form factor phase as a function of the $\pi\pi$ invariant mass. The solid curve corresponds to the central value (neglecting inelastic contributions), and the blue dashed ones depict one standard deviation uncertainties, assuming $|\delta_{+}(s)-\delta_{T}(s)|=2\delta_{+}^{inel}(s)$. For comparison the central value of $\delta_{+}(s)$, according to ref. [75] (see fig. 2), is shown by a black dashed line.

### v2

@@ Abstract
We apply the effective field theory formalism that was used to study CP violation induced by heavy new physics in the $\tau\to K_{S}\pi\nu_{\tau}$ decays to the other two-meson tau decay channels. We focus on the rate and the forward-backward asymmetries, that are predicted using current bounds on the complex Wilson coefficients of the effective Lagrangian. We discuss our outcomes for the modes with $(K/\pi)\pi^{0}$ and $KK_{S}$, that can be studied at Belle-II and a super-tau-charm facility. Our main finding is that current and forthcoming experiments would be sensitive to the maximum allowed CP rate asymmetry in the $K^{\pm}K_{S}$ modes if a precision of $5\%$ is reached on this observable, that can check as well the BaBar anomaly in $K_{S}\pi$. For the $\pi^{\pm}\pi^{0}$ channels, new physics would be difficult to probe at present. Disentangling new sources of CP violation would be most challenging in $K^{\pm}\pi^{0}$ and the other modes.
arxiv: 2409.05588
@@ 1 Introduction
Violation of the invariance under the combined symmetry operations of charge conjugation (C) and parity (P), CP, is a required ingredient -within any Lorentz invariant local quantum field theory with a Hermitian Hamiltonian- to understand the baryon asymmetry of the Universe, as pointed out by Zakharov in his seminal paper Sakharov:1967dj . It also needs baryon number violation (which could be triggered violating initially the lepton number Fukugita:1986hr ), non-conservation of C, and interactions out of thermal equilibrium. In principle all these conditions can be met within the Standard Model (SM) Glashow:1961tr ; Weinberg:1967tq ; Salam:1968rm , though it fails to explain the current matter-antimatter asymmetry ParticleDataGroup:2024cfk . This conundrum enhances the case for searches of additional sources of CP violation, that can eventually solve this puzzle.
@@
In the quark sector, CP violation in the SM arises only from an irreducible phase in the Cabibbo-Kobayashi-Maskawa (CKM) mixing matrix Cabibbo:1963yz ; Kobayashi:1973fv , and has been measured in $K$, $B_{(s)}$ and $D$ mesons Christenson:1964fg ; NA31:1988eyf ; KTeV:1999kad ; NA48:1999szy ; BaBar:2001pki ; Belle:2001zzw ; BaBar:2004gyj ; Belle:2004nch ; LHCb:2012xkq ; LHCb:2013syl ; LHCb:2019hro . To date, no data from hadron decays or meson mixing seems to require CP violation beyond the SM ParticleDataGroup:2024cfk . Conversely, CP violation in the lepton sector has not been discovered yet, although there are hints for it both in neutrino oscillation measurements by Nova NOvA:2021nfi and T2K T2K:2023smv , and in global fits ParticleDataGroup:2024cfk . If confirmed, it would remain to be seen if it can be explained by the single CP violating phase that appears in the lepton analogue of the CKM matrix, the Pontecorvo-Maki-Nakagawa-Sakata matrix Maki:1962mu ; Pontecorvo:1967fh ; Katayama:1962mx , or new physics is called for.
@@
Hadron tau decays are not only a privileged arena for learning about non-perturbative QCD in a clean environment, but also a very useful tool in the quest for new physics Pich:2013lsa ; Davier:2005xq . Regarding CP violation, this was emphasized for the first generation B-factories, BaBar and Belle, and forthcoming facilities Tsai:1996ps ; Kuhn:1996dv ; Bigi:2012km ; Bigi:2012kz ; Kiers:2012fy . CP non-conservation in semileptonic tau decays was boosted by the anomaly reported by BaBar BaBar:2011pij on the difference between the decay rates of a $\tau^{+}$ and a $\tau^{-}$ into final states with a correspondingly charged pion and a $K_{S}$ (inclusive in possible neutral pions), normalized by their sum. While this CP violating observable should be $(0.36\pm 0.01)\%$ in the Standard Model –due to neutral kaon mixing and accounting for the precise BaBar experimental setting– Bigi:2005ts ; Calderon:2007rg ; Grossman:2011zk , the measurement came out with opposite sign, $-(0.36\pm 0.23\pm 0.11)\%$ (with statistical and systematic errors in turn), being $2.8$ standard deviations away from the theory prediction. Although there were claims that this effect could be due to heavy new physics (particularly of tensor type Devi:2013gya ; Dhargyal:2016kwp ; Dighe:2019odu ), the key equality of the tensor and vector form factors phases in the elastic region allowed to disfavor this explanation unless an unnatural fine-tuning was involved Cirigliano:2017tqn (see also refs. Rendon:2019awg ; Chen:2019vbr ; Chen:2020uxi ; Sang:2020ksa ; Chen:2021udz ).  [^1] No CP violation was found in the binned analysis of these decays by Belle Belle:2011sna . However, its precision was not enough to clarify the BaBar anomaly (neither in the pioneering CLEO studies CLEO:1998lwx ; CLEO:2001lhp ).
@@
The aim of this paper is to extend the work done in the previous references to the other two-meson tau decays of interest ($\pi\pi$, $K\bar{K}$ and $K\pi^{0}$ channels).  [^2] The $\pi\eta^{(\prime)}$ channels have not been measured yet, and the hadronic uncertainty in them is big Descotes-Genon:2014tla ; Escribano:2016ntp , so any current prediction would be subject to large uncertainties. CP conserving new physics imprints in these decays were the subject of ref. Garces:2017jpz . We will comment briefly on the $K(\eta/\eta^{\prime})$ channels, also subject to large uncertainties, at the end of section 5.  [^3] Model-dependent studies have been carried out in e. g. refs. Delepine:2006fv ; Delepine:2007qg ; LopezCastro:2009da ; Kimura:2012bwp ; Gao:2012su ; Delepine:2018amd ; Faisel:2023ldd .  [^4] We will not address in this work the case of polarized taus, which would allow further tests of $CP$ and time-reversal symmetry violations, along the lines of ref. Chen:2022nxm . In the same way that heavy new physics contributions to the CP violating $\tau\to K_{S}\pi\nu_{\tau}$ observables is very much restricted, model-independently, upon translating the constraints on the SM effective-field theory (EFT) Buchmuller:1985jz ; Grzadkowski:2010es operators to the low-energy EFT (LEFT) ones, it will happen similarly in the other two-meson tau decay modes. Our purpose is to compute the maximum amount of CP violation that can be due to heavy new physics in these processes. In this way, any observation beyond our predicted upper limits  [^5] Within the Standard Model CP violation is negligibly small in all tau decays without a neutral Kaon, see for instance ref. Delepine:2005tw . will either be due to light new physics or signal underestimated uncertainties, which would deserve a close scrutiny. The discussed physics shall be of interest for Belle-II Belle-II:2018jsg and future super-tau-charm factory Achasov:2023gey ; Achasov:2024eua ; CPVatSTCF searches.
@@ 2 Low-energy Effective Field Theory Computation
The most general Fermi-like effective Lagrangian for the $\tau^{-}\to\bar{u}D\nu_{\tau}\,(D=d,s)$ decays is given by Cirigliano:2009wk ; Garces:2017jpz ; Cirigliano:2017tqn   [^6] In our conventions, $\sigma^{\mu\nu}=\frac{i}{2}[\gamma^{\mu},\gamma^{\nu}]$ and $\epsilon^{0123}=+1$.
@@
The other decay channels of interest are trivially obtained from the $K^{-}\pi^{0}$ one, replacing $p_{K}$ ($p_{\pi}$) by the charged (neutral) meson momentum and substituting the $1/2$ factor in eq. (2) by $1/\sqrt{2}$ for the $K^{-}K^{0}$ and $\bar{K}^{0}\pi^{-}$ modes and by $1$ in the $\pi^{-}\pi^{0}$ channel. The $K^{-}\eta^{(\prime)}$ channels are a bit more complicated. Relative to the $K^{-}\pi^{0}$ modes, one needs to substitute $1\to c^{V}_{K^{-}\eta^{(\prime)}}$ in the $F_{+}(s)$ term and $1\to-c^{S}_{K^{-}\eta^{(\prime)}}$ in the $F_{0}(s)$ term, where $c^{V}_{K^{-}\eta^{(\prime)}}=-\sqrt{\frac{3}{2}}$, $c^{S}_{K^{-}\eta}=-\frac{1}{\sqrt{6}}$ and $c^{S}_{K^{-}\eta^{\prime}}=\frac{2}{\sqrt{3}}$ Escribano:2013bca . The scalar matrix element, $H$, demands to modify $1\to c^{S}_{K^{-}\eta^{(\prime)}}$, in order to obtain the $K^{-}\eta^{(\prime)}$ result from the $K^{-}\pi^{0}$ expression. For the tensor matrix element $H^{\mu\nu}$, the $K^{-}\eta^{(\prime)}$ result is found from the $K^{-}\pi^{0}$ one by changing $1\to c^{V}_{K^{-}\eta^{(\prime)}}$.
The observables of interest are quoted in the following. All of them check the results in ref. Chen:2021udz accounting for the relative factor of $1/2$ entering our results for the (differential) decay width:
@@
where $\lambda(a,b,c)=a^{2}+b^{2}+c^{2}-2(ab+ac+bc)$ and $S_{EW}\sim 1.02$ Marciano:1988vm encodes the universal short-distance electroweak radiative corrections (which only affect the SM contribution, but are written here as an overall factor for simplicity, since the induced error of this approximation is subleading).  [^8] We are not considering observables inclusive in photon emission, so that the process-dependent long-distance electromagnetic correction Cirigliano:2001er ; Cirigliano:2002pv ; Miranda:2020wdg is not applied. Integrating the angular dependence,  [^9] We recall that $\alpha$ is the angle between one of the mesons momentum and the tau in the hadronic rest frame Kuhn:1992nz ; Belle-II:2018jsg . we get
@@
where $\delta_{+,T}(s)$ denote the phases of the vector and tensor form factors, in turn. The expression in the second and third lines of eq. (11) corresponds to the new physics (NP) contribution to $A_{CP}^{rate}$, which coincides with $A_{CP}^{rate}$ for those modes without neutral Kaons. In the $\tau^{\pm}\to K^{\pm}K_{S}\nu_{\tau}$ case, this observable is dominated by the SM contribution, according to Devi:2013gya
@@
For completeness, we note that these observables can be made bin-dependent straightforwardly, following ref. Chen:2021udz , namely
@@ 3 Form-factors input
We will use the dispersive form factors that were employed in e. g. ref. Gonzalez-Solis:2020jlh , which bound the modulus of the Wilson coefficients $\hat{\epsilon}_{S,T}^{D}$ using exclusive semileptonic tau decays into one and two mesons. We will illustrate this hadron input here, taking as an example the simplest di-pion $F_{+}$ case, Pich:2001pj ; GomezDumm:2013sib ; Gonzalez-Solis:2019iod . The other dispersive form factors in this channel and those for the remaining decay modes were worked out in e. g. refs. Descotes-Genon:2014tla ; Miranda:2018cpf ; Jamin:2000wn ; Jamin:2001zq ; Moussallam:2007qc ; Boito:2008fq ; Boito:2010me ; Antonelli:2013usa ; Bernard:2013jxa ; Rendon:2019awg ; Escribano:2013bca ; Escribano:2014joa ; Gonzalez-Solis:2019lze and agree well with data, given the present uncertainties. We will briefly comment on the sources of theory uncertainty for $F_{+}^{\pi\pi}$, which is the best known, below.
@@
where $\alpha_{1,2}$ are subtraction constants (the other one is fixed by the conservation of the vector current, so that $F_{+}(0)=1$) giving the slope and curvature of the threshold expansion of $F_{+}(s)$, where Chiral Perturbation Theory provides useful constraints Weinberg:1978kz ; Gasser:1983yg ; Gasser:1984gg ; Cata:2007ns . In practice, since the phaseshift $\delta_{+}(s)$ is not known up to infinity, several approaches are used to account for the associated uncertainty, as detailed in the quoted refs. In the resonance region we obtain the seed for $\tan\delta_{+}(s)=\frac{\Im mF_{+}(s)}{\Re eF_{+}(s)}$ from Resonance Chiral Theory ($R\chi T$Ecker:1988te ; Ecker:1989yg , with parameters encoding the resonance properties (their pole positions and interferences among them) that are fitted to data. Our form factors comply with the leading asymptotic behaviour predicted by perturbative QCD Lepage:1980fj .
@@
In the remainder of this section, we summarize the main sources of theory uncertainty, analyzed minutely in ref. Gonzalez-Solis:2019iod for $F_{+}^{\pi\pi}(s)$. In the elastic zone (which, neglecting the phase-space suppressed four-pion contributions extends up to the di-Kaon channel, with invariant mass around a GeV), the form factor phase is known with extreme precision, thanks to the Roy-Steiner equation analyses of pion scattering data by the Madrid Garcia-Martin:2011iqs and Bern Caprini:2011ky groups (and to the Watson theorem Watson:1952ji that ensures the equality of the scattering and form factor phases in this region). Ref. Gonzalez-Solis:2019iod , working within the excellent approximation of isospin symmetry, uses these results in the elastic domain, including the possible small variation of the precise energy at which the phase reaches $\pi/2$, signaling the $\rho(770)$ resonance position. Another source of uncertainty corresponds to the matching point between the elastic phase and the one obtained using as seed the $R\chi T$ expression, where values in the interval $[0.9,1]$ GeV are employed. For the $\rho^{\prime(\prime)}$ resonances widths, both the simplest case with only the di-pion channel and also that adding the di-Kaon and $\omega$-$\pi$ cuts were considered. As commented before, there is a lack of experimental information on $\delta_{+}^{\pi\pi}(s)$ above a given energy, and we only know that it must approach $\pi$ asymptotically Leutwyler:2002hm . Several methods quantified this uncertainty in ref. Gonzalez-Solis:2019iod : cutting the upper limit ($s_{cut}$) of the dispersive integral in $4$, or $9$ GeV2, or keeping it up to $\infty$; modifying the phase for $s\sim s_{cut}$ to avoid spurious singularities; including additional inelastic contributions via a conformal polynomial, etc. It was also verified that the number of subtraction constants was optimal using the reduced $\chi^{2}$ as a criterion. The impact of the $\tau^{-}\to K^{-}K_{S}\nu_{\tau}$ data was also explored, although it is not yet precise enough to improve our knowledge of the $\rho^{\prime}-\rho^{\prime\prime}$ interference.
@@
In this way, ref. Gonzalez-Solis:2019iod quantified the uncertainty of $F_{+}^{\pi\pi}(s)$, whose modulus and phase are represented in figures 1 and 2. Following ref. Cirigliano:2017tqn , we estimated the difference between the vector and tensor form factors in the inelastic region  [^10] They coincide in the elastic one, as elegantly explained in ref. Cirigliano:2017tqn . ($inel$) using $\delta_{+}(s)-\delta_{T}(s)=2\delta_{+}^{inel}(s)$. The tensor form factor phase, $\delta_{T}(s)$, thus obtained is compared in figure 3 to $\delta_{+}(s)$. Figure 4 plots $|F_{T}(s)|/F_{T}(0)$. In this case, we lack information on the slope parameters, so we use a once-subtracted dispersion relation, which increases the sensitivity to $s_{cut}$. As a result, the uncertainty on $|F_{T}(s)|$ is larger than for $|F_{+}(s)|$, despite being their phases identical in the elastic regime.
@@
Figure 1: Pion vector form factor modulus as a function of the $\pi\pi$ invariant mass. The solid curve represents the central value, and the dashed lines cover the one sigma uncertainties, according to ref. Gonzalez-Solis:2019iod .
@@
Figure 2: Pion vector form factor phase as a function of the $\pi\pi$ invariant mass. The solid curve represents the central value, and the dashed lines cover the one sigma uncertainties, according to ref. Gonzalez-Solis:2019iod .
@@
Figure 3: Pion tensor form factor phase as a function of the $\pi\pi$ invariant mass, assuming $|\delta_{+}(s)-\delta_{T}(s)|=2\delta_{+}^{inel}(s)$. The solid curve corresponds to the central value (neglecting inelastic contributions), and the black dashed inner line depicts $\delta_{+}(s)$ for comparison, according to ref. Gonzalez-Solis:2019iod (see figure 2).